Please show all of your work! Thank you[Problem 7-63] If 8% is considered the minimum attractive rate of return, which alternative should be selected using an incre
Answer #1

MARR = 8%

Calculating the incremental cash flow between X – Y

EOY

X

Y

X – Y

0

-5000

-5000

0

1

-3000

2000

-5000

2

4000

2000

2000

3

4000

2000

2000

4

4000

2000

2000

Calculating the Incremental IRR of X – Y using the Trial and Error Method.

Calculate the Present Worth of the incremental cash flow at MARR of 8%

PW = -5,000 (P/F, 8%, 1) + 2,000 (P/A, 8%, 3) (P/F, 8%, 1)

PW = -5,000 (0.92593) + 2,000 (2.57710) (0.92593) = 142.78

The present worth is positive. Increase the MARR to 10% to get negative PW. Calculate the PW using interest rate = 10%

PW = -5,000 (P/F, 10%, 1) + 2,000 (P/A, 10%, 3) (P/F, 10%, 1)

PW = -5,000 (0.90909) + 2,000 (2.48685) (0.90909) = -23.90

Using interpolation to calculate the Incremental IRR

IRR = 8% + [142.78 – 0 ÷ 142.78 – (-23.90)]*2% = 9.7%

IRR of Incremental Cash flow (X – Y) > MARR, Select Alternative X